For the reaction: 2N205 (g) →4NO2(g) +O2(g) The experimental data for this reaction is t:0.0 s [N205] = 0.80 M NO2):0.0 M [O21:00 M t; 20.0 s[N205]=0.20 M INO 1.2M [O2-0.30 M was the average rate of formation of NO2 and O2 during the 20 second time interval? The average rate of formation of NO2 was +0.030 M/s, and the average rate of formation of 02 was +0.030 Ms The average rate of formation of NO2 was +0.030 M/s, and the average rate of formation of 02 was +0.015 Ms The average rate of formation of NO2 was +0.060 M/s, and the average rate of formation of O2 was +0.0075 Ms. The average rate of formation of NO2 was +0.060 M/s, and the average rate of formation of 02 was +0.015 M/s. The average rate of formation of NO2 was +0.060 M/s, and the average rate of formation of O2 was +0.030 Ms
Answer #1

average rate of formation of NO2 = [NO2] finall - [NO2] initial / ( time change)

= ( 1.2- 0)M /( 20 s) = 0.06 M/s

average rate of formation of O2 = [O2] final [ O2] initial / ( time change)  

= ( 0.3-0)M / 20s = 0.015 M/s

Thus answer is option 4